Technical Observing AGW1 & AGW2 Testing w/LUCI1 & LUCI2 20131212 UT

Observers: JHill, DMiller (Tucson), DThompson (LBTO)
Telescope Operator: GBechetti (LBTO)
AO Support: DMiller (Tucson)
Software Support: MDeLaPena (Tucson)
Telescope Support: MMidkiff (LBTO)
Instrument Support: ESolheid (LBTO)
Engineering Support: DAshby (Tucson)

Summary

Opened late because of humidity.

Worked on AGW1 and AGW2 probe alignment and wfs alignment, but closed again at 04:35 UT for humidity before we got to transformations. Found and fixed the critical problem (InformationTechnology 4960) where AGW2 configuration was not read correctly since ~ 8-Nov.

Details

Open momentarily, then close for humidity.

Open

01:35 Open again

01:43 sync ACQUIRE presets with LUCI/LUCI (from IRAF). Both sides are off in pointing about the same amount.
SX IE=-45 CA=-26
DX IE=-12 CA=-185

RFBG lenslet positions

01:49 sync ACTIVE presets to BS9101

01:50 Doug is fixing the DX lenslet positions, and the guiding hotspot positions - Why did they move?

Testing Issue 4514

John does this test while Doug is working on the hotspot data. Binocular alignment really takes two people.

Still with the LUCI/LUCI ACTIVE preset running.......guiding and wfs corrections are blocked at GCS, but still being saved.

02:03 Move SX pointing with: ptincrement 0 IE 40; ptincrement 0 IE -40; ptincrement 0 CA -40; ptincrement 0 CA 40

Then repeat moving the pointing on the DX side.

02:10 Test complete -- with no trace of Issue 4514 down to the level of the image motion in ~1 arcsec seeing. Issue can be closed.

What else could cause the lenslet alignment to change?

02:13 Init DX M3 to make sure it is in the correct position. It was, star image returned within a couple arcsec.

Issue 4961 which wasn't an Issue

02:19 An async preset on the right side caused the left side to lose the guide star.

02:20 Resend the left async preset which did not move the right guide star.

Doug adjusted the DX AGW2 guide hotspot to 269,265 pixels.

02:32 async ACQUIRE preset on DX moved the SX star away.

02:32 async ACQUIRE preset on SX

The next morning, Michele found that the above 2 DX presets were sent with a CLEARPRESET Binocular Control Flag (BC=2 in IRAF speak). This does not protect the SX target, so everything worked correctly. "async" presets from IRAF should have BC=INDEF so as not to send any Binocular Control Flag.

AGW1 and AGW2 Guide Probe Centering

Following the procedure in BGTechProcedures, but with the extra complication of Doug and John doing it in parallel on the two sides. It worked pretty well.

Data for the right in /home/lbto/data/20131212_agw2/ and I'm not sure where Doug put the left data.

02:40 DX x=-0.22 y=611.8 EL=64.5 deg
right_guide 0001 center of circle at 122.5, 95.5

02:44 DX x=0.00 y=612.5 or encoder coordinates 0.0,1001002.000
right_guide 0002 center of circle at 126,105
x2 plus 17,15 gives unbinned coordinates of 269,225 ---> so need dY of +40 pixels to match hotspot
40 pixels = 2 arcsec = 1200 microns.
Subtract 1200 microns from AGW2 : RAD : ooffset to move up in Y. 419522 ---> 418322.

03:0x SX x=0.2 y=612.2
left_guide 0001 center of circle at 131.4,109.3

03:07 SX x=0.0 y=612.5
left_guide 0002 center of circle at 129.1, 112.5
correct binning 258.2, 225.0
add offsets 273.2,264.0
corrections: -18,-34 pixels
corrections: +0.9,+1.7 arcsec
corrections: +540,+1020 microns
corrections: +36 counts, --

Change AGW1 : ROT : homeoffset from 35548 ---> 35584
Change AGW1 : RAD : ooffset from 417855 ---> 418875

03:21 New GCS v13 installed "with protection around sourceextractor"

03:4x Stop GCSs, Stop AGws, stop oacserver, start oacserver, start AGws, start GCSs

03:46 sync ACQUIRE presets on both sides
SX IE=-71 CA=-17
DX IE=-2 CA=-194

03:48 sync ACTIVE presets on both sides
Stop guiding
ptincrement 2 CA 5
slew both rotators to 0 deg

03:53 Two almost simultaneous 90 sec circle exposures on the two sides.
right_guide 0003 and left_guide 0003

Move both probes to x=0.0 y=612.5

03:56

right_guide 0004 has the circle at 125,104.5 --- InformationTechnology DIDN'T MOVE from our previous adjustment!
so make another adjustment of AGW2 : RAD : ooffset to 417122

left_guide 0004 has the circle at 121,98 translating to 257,235 --- OK

Debugging Issue 4960 - Why doesn't the AGW2 probe move when we change the configuration?

04:12 Stop GCSs, Stop AGws, stop oacserver, start oacserver, start AGws, start GCSs

04:18 ACTIVE preset

setxy -u 2 -x 0.0 -y 612.5

right_guide 0005 has the circle at 125,105 (unchanged)

04:28 Removed a stray space after the RAD offset line for AGW2

04:30 Stop GCSs, Stop AGws, stop oacserver, start oacserver, start AGws, start GCSs

Removing the space didn't help.

Close

04:37 Close for rising humidity, continue working on Issue 4960 with dome closed.

04:40 Copy oacontrol.conf.20131025 into oacontrol.conf (old version from before the AGW7/AGW8 additions for PEPSI)

04:42 Stop GCSs, Stop AGws, stop oacserver, start oacserver, start AGws, start GCSs

Check positions with getxy -u 1 -o 16 and getxy -u 2 -o 16
AGW1 0,612.5 ---> 0 1006955.000
AGW2 0,612.5 ---> 0 1001001.000

04:52 edit RAD ooffset for both AGW1 and AGW2 in the "old" 20131015 copy of oacontrol.conf now installed.

04:53 Stop GCSs, Stop AGws, stop oacserver, start oacserver, start AGws, start GCSs

AGW1 0,612.5 ---> 0 1001648.000
AGW2 0,612.5 ---> 0 1001001.000

AGW1 changed, but AGW2 did not.

05:02 Stop GCSs, Stop AGws, warmboot agw-control, start AGws, start GCSs

05:45 Change oacontrol.conf.default from oacontrol.conf.default.20131115.agw123478.PEPSI to oacontrol.conf.default.20131107

BUT that one still has duplicate RAD ooffset entries for AGW2

cp -p oacontrol.conf.20131212 oacontrol.conf (to put back our version from earlier tonight)

05:46 Stop GCSs, Stop AGws, stop oacserver, start oacserver, start AGws, start GCSs

06:25 Finished another restart of everything to recover from AzCam issues.

06:46 cp -p oacontrol.conf.default.20131115.agw123478.PEPSI oacontrol.conf.default, and then commented out the duplicate entry for AGW2 : RAD : ooffset

06:51 Stop GCSs, Stop AGws, stop oacserver, start oacserver, start AGws, start GCSs

AGW2 probe now goes to a different value. 0,612.5 ---> 0.0,1004525.000

07:00 returned AGW2 : RAD : ooffset to its previous (20131025) value of 419522.

07:02 Stop GCSs, Stop AGws, stop oacserver, start oacserver, start AGws, start GCSs

AGW2 probe now goes to a different value. 0,612.5 ---> 0.0,998284.000

Thus, the problem (InformationTechnology 4960) is solved, and we can once again adjust the AGW2 probe center position. Probably now the 20131025 settings for AGW2 are in use and good - Doug will need to put the IRTC-based AGW2 tranform back in.

07:30 Geno plays "Goodnight Sweetheart" as we decide that the weather is hopeless (humidity, wind, visibility < 50 meters).

-- JohnHill - 12 Dec 2013
Topic revision: r5 - 25 Jul 2014, DouglasFisher
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback